OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5.

OpenStudy (anonymous):

Parallel---> Same slope---->m=-1/2 Then use this: \[y-y_1=m(x-x_1)\] Where (x1, y1)=(4,-6) Tell us what do you get

OpenStudy (anonymous):

x - 8

OpenStudy (anonymous):

y+6=(-1/2)*(x-4) Start from there.

OpenStudy (anonymous):

y = -2x + 2

OpenStudy (anonymous):

\[y+6=\frac{-1}{2}(x-4)=\frac{-x}{2}+2\] And now?

OpenStudy (anonymous):

y = -1/2x - 4

OpenStudy (anonymous):

You got it! Congratulations!

OpenStudy (anonymous):

PD:Thanks for the "Best responnse" ;)

OpenStudy (anonymous):

:D